how do you find the slope of a line paralell to the graph of each eqution?
1: 9x -5y = 4
2: -x + 3y = 6
3: 6x - 7y = 10
4: x= -4
5: -3x - 5y = 6

Answer by stanbon(75887)

how do you find the slope of a line paralell to the graph of each eqution?
1: 9x -5y = 4
y = (9/5)x - (4/5)
slope = 9/5
------------------------
2: -x + 3y = 6
y = (1/3)x + 2
slope = 1/3
-----------------------
3: 6x - 7y = 10
y = (6/7)x - (10/7)
slope = 6/7
------
4: x= -4
slope is "undefined"
-----------------------
5: -3x - 5y = 6
y = (-3/5)x - (6/5)
slope = -3/5
=====================
Cheers,
Stan H.
=====================